which of the following statement is true:
1. a number divisible by 6 is also divisible by 3
2. a number divisible by 3 is also divisible by 9
3. a number divisible by 5 is also divisible by 10
4. a number divisible by 3 is also divisible by 6

Answer by funmath(2933)

1. a number divisible by 6 is also divisible by 3
Is true because the divisibility rule for 6 is if the number is even and divisible by 3 then the number is divisible by 6. Being divisible by three is part of the rule.
Observe:
12/6=2 and 12/3=4
24/6=4 and 24/3=8
ect. ect. ect.
Happy Calculating!!!